# KNN-MMD: Cross Domain Wi-Fi Sensing Based on Local Distribution Alignment

### v1

@@ Abstract
As a key technology in Integrated Sensing and Communications (ISAC), Wi-Fi sensing has gained widespread application in various settings such as homes, offices, and public spaces. By analyzing the patterns of Channel State Information (CSI), we can obtain information about people’s actions for tasks like person identification, gesture recognition, and fall detection. However, the CSI is heavily influenced by the environment, such that even minor environmental changes can significantly alter the CSI patterns. This will cause the performance deterioration and even failure when applying the Wi-Fi sensing model trained in one environment to another. To address this problem, we introduce a K-Nearest Neighbors Maximum Mean Discrepancy (KNN-MMD) model, a few-shot method for cross-domain Wi-Fi sensing. We propose a local distribution alignment method within each category, which outperforms traditional Domain Adaptation (DA) methods based on global alignment. Besides, our method can determine when to stop training, which cannot be realized by most DA methods. As a result, our method is more stable and can be better used in practice. The effectiveness of our method are evaluated in several cross-domain Wi-Fi sensing tasks, including gesture recognition, person identification, fall detection, and action recognition, using both a public dataset and a self-collected dataset. In one-shot scenario, our method achieves accuracy of 93.26%, 81.84%, 77.62%, and 75.30% in the four tasks respectively. To facilitate future research, we will make our code and dataset publicly available upon publication.
@@ I Introduction
Existing prominent progress in Deep Learning (DL), wireless communication, and sensing have envisioned that the future networks will not only serve the sole purpose of data delivery but also actively and cooperatively collect data to recognize the environment. The Integrated Sensing and Communication (ISAC) devices are expected to be massively deployed in the real world, supporting and providing a large variety of tactile intelligent services and applications.
One materialization of the concept of ISAC is Wi-Fi sensing, where the Wi-Fi devices serve as the radar sensors in local areas [1, 2]. The advantages of Wi-Fi sensing compared with other sensing method are given in Table I. First, Wi-Fi sensing is contactless as the object to be detected does not need to carry a device like a smartphone or smartwatch. Second, Wi-Fi sensing is privacy-preserving compared with camera. Third, Wi-Fi sensing is insensitive to occlusions and illumination conditions since its electromagnetic wave operates on a frequency band that can penetrate walls. Moreover, Wi-Fi sensing is ubiquitous, with over $19.5$ billion Wi-Fi devices deployed globally [3]. With an appropriate algorithm that enables its sensing function, ubiquitous sensing function can be realized in various environments, such as homes, offices, and public spaces. Last but not least, Wi-Fi sensing has lower cost compared to dedicated radars.
@@
Given the released tool that works with some specific types of commercial Network Interface Cards (NIC) [4], the Channel State Information (CSI) can be obtained, which describes how the wireless signals fade during the propagation from the sender to the receiver. By using this CSI data collection tool, a series of research employing commercial Wi-Fi NICs for different kinds of tasks. Moreover, a dedicated IEEE standard for Wi-Fi sensing named IEEE 802.11bf is approved [5] to enable faster development and practicalization of Wi-Fi sensing technologies. However, since the Orthogonal Frequency Division Multiplexing (OFDM) waveforms employed by Wi-Fi are designed for higher communication rate, it is hard to use simple physical models to extract information about what happened in the real world from the CSI measurements.
As a result, most Wi-Fi sensing methods rely on machine learning, especially for fine-grained tasks like gesture recognition [6]. However, even minor environmental changes can drastically alter the CSI patterns and significantly degrade the performance of the trained model. This environmental variability leads to a domain shift - a discrepancy between the training data distribution and real-world deployment conditions. Given the strong dependence of the training data, such domain shifts can severely undermine the generalization capability. This scenario is known as a cross-domain problem in machine learning, where the source domain (training data) and target domain (testing data) have different underlying distributions or feature spaces.
@@
As the most common method for cross-domain tasks, Domain Adaptation (DA) has been widely used in cross-domain Wi-Fi sensing. DA focuses on leveraging knowledge from a source domain to improve model performance on a related but different target domain. In cross-domain Wi-Fi sensing, the most commonly used methods are Few-Shot Learning (FSL) methods and Domain Alignment (DAL) methods:
  • $\bullet$
    FSL methods involve training a network by using abundant data in the source domain and limited data in the target domain. For instance, the Deep Similarity Evaluation Network [8, 9], Matching Networks [10, 11], and Siamese Network [12, 13] are all designed or applied to learn a transferable similarity evaluation ability from the training dataset and apply this learned knowledge to new testing domains. Prototypical Network [14, 15] aims to learn an embedding function that maps samples from the same class close to a prototype in the embedding space, while samples from different classes are well separated. However, the model performance on the testing set is not stable during training (as referenced in [16] and Section IV-D). This instability in the model performance makes it difficult to use in practical scenarios.
  • $\bullet$
    DAL methods aim to map the source and target domains to the same distribution. Among these, the traditional and straightforward methods, such as Maximum Mean Discrepancy (MMD) [17] and its improvement Multiple Kernel Maximum Mean Discrepancy (MK-MMD) [18], are the most widely used in Wi-Fi sensing [16, 19, 20, 21, 22, 23, 24, 25]. However, as shown in Fig. 1, the traditional DAL methods do not guarantee model performance (as referenced in [26] and Section III-B).
To overcome these problems of DA methods in cross-domain Wi-Fi sensing, we propose a novel FSL framework, called K-Nearest Neighbors Maximum Mean Discrepancy (KNN-MMD). In our method, we re-model the cross-domain task and propose a local distribution alignment method to improve the traditional DAL approach. Additionally, our method ensures that the support set does not participate in the training process of the network, allowing us to apply an early stop strategy based on the accuracy in the support set. This enables our training process to stop at a point where the performance in the target domain is always good. In summary, the main contributions of this work are:
(1) A FSL Framework based on Local Alignment: A FSL framework called KNN-MMD is designed for cross-domain Wi-Fi sensing. As the global alignment strategy of traditional MMD cannot guarantee model performance, a more effective approach has been proposed that aligns the data distributions from the source and target domains within each category. We first use KNN to generate pseudo labels for the target domain and generate a high-confidence help set. Then the help set and training set are used for local alignment during training classification network on training set.
(2) An Early Stop Strategy based on Support Set: As the support set does not participate in training in our proposed method, we can determine when to stop training using an early stop strategy on the accuracy and loss in the support set, which is difficult to incorporate in traditional domain adaptation methods. This makes our approach more practical.
@@
(3) Experiment Evaluation: Experimental results show the superior performance of our method in multiple Wi-Fi sensing tasks, including gesture recognition, people identification, fall detection, and action recognition. Our method outperforms most models in both performance and stability.
@@
II Problem Description and Benchmark Solutions
@@ II-A Wi-Fi Sensing Principle
II-B Cross Domain Wi-Fi Sensing
@@
In this paper, we consider the WiGesture dataset [27], which contains the CSI collected from 8 individuals performing 6 different gesture actions and a self-collected WiFall dataset, including CSI collected from 10 individuals performing 5 different actions. For the gesture recognition or action classification task, we take different people as different domains. And for the people identification task, we take different actions as different domains.
In the following of this paper, the variables $x$ and $y$ refer to the input data and output label, respectively. $\theta$ represents the network parameters. $P_{s}$ and $P_{t}$ represent the probabilities in the source domain and target domain, respectively. Furthermore, the focus of this paper is primarily on the few-shot cross-domain task. Most available data is in the source domain expect $n$ labeled samples for each category in the target domain. This scenario is often referred to as the n-shot problem. In addition, we also have some unlabeled data from the target domain. Our objective is to accurately classify these unlabeled data. For the sake of simplicity, we will refer to the labeled data from the source domain as the training set, and the labeled data from the target domain as the support set.
Next, we use WiGesture dataset to illustrate the domain shift in Wi-Fi sensing. Fig. 2 depicts the distribution of mean and standard deviation of CSI amplitude across various individuals and actions. It is evident that even for the same action, individuals exhibit distinct data distributions. Likewise, Fig. 3a and 3b display the Uniform Manifold Approximation and Projection (UMAP) based data reduction [28] results (we will introduce this data dimension reduction method in Section III-D1). It is apparent that within a single individual, different actions exhibit significant dissimilarities. However, when considering different individuals, the boundaries separating different actions become less distinct. These findings imply that it is challenging to generalize the learned features from one set of individuals to others. In Fig. 3c, a Resnet18 [29] is trained by setting the people with IDs 1-7 as the source domain and people with ID 0 as the target domain. It can be observed that the trained Resnet18 can distinguish different actions in the source domain but fail to distinguish different actions in the target domain. This is mainly because even for the same action, different people have distinct behavioral patterns, which can lead to significant gaps in the CSI patterns between individuals. As a result, to make the model perform well in the target domain, it is necessary to address the domain shift (cross-domain) problem.
@@
II-C Domain Adaptation Methods
@@
Metric-based Method Learning-based Method Domain Alignment Method Our Method
Representative Methods KNN[30], K-means[31] Siamese[12], Triplet Network[32] MMD[17], GFK[33, 34] KNN-MMD
Sensitivity to Quality of Support Set High Moderate None Low
Stability Low Low Low High
Assumption $P_{t}(y|x)=P_{s}(y|x)$ No Some methods require it[35] Yes No
As the most common method in cross-domain tasks, DA can be mainly classified into three types: metric-based methods, learning-based methods, and DAL methods. However, each of these approaches has its own shortcomings, as shown in Table II.
@@
DAL methods aim to map the source domain and target domain to a common feature distribution, where the data in the two domains can have the same distribution in the new space. However, most DAL methods, such as MMD [17], Kernel Mean Matching (KMM) [39], and Subspace Geodesic Flow (SGF) [40], rely on the assumption that the conditional probabilities in the source domain and target domain are the same, i.e., $P_{t}(y|x)=P_{s}(y|x)=P(y|x)$. To estimate $P(y|x)$, the network parameters $\theta$ need to be learned to achieve $P(y|x;\theta)=P(y|x)$. However, due to $P_{t}(x)\neq P_{s}(x)$, the $P(y|x;\theta)$ learned from the source domain cannot generalize to the target domain. Consequently, these traditional methods can be seen as finding an intermediate space where the input data from the source domain and target domain are mapped to the same distribution, i.e., $P_{t}(x)=P_{s}(x)$. This allows the model to learn the parameter $\theta$ more effectively and accurately. However, in many practical scenarios, the assumption $P_{t}(y|x)=P_{s}(y|x)$ is not met. This has been demonstrated in our experiments (Section IV-C) as well as our previous research [16], where traditional DAL methods even failed in cross-domain Wi-Fi sensing tasks.
@@ III-C Model Structure
In our model, as shown in Fig. 5, we utilize Resnet18[29] as the feature extraction component and Multilayer Perceptrons (MLP) as the classifier. Specifically, we begin with some normalization operations, which we have observed to be essential for the success of the model.
@@ IV-C Experiment Result
Method Scenario Gesture Recognition People Identification Fall Detection Action Recognition Average Accuracy
Resnet18[29] in-domain 80.75% 86.75% 91.88% 70.50% 82.47%
zero-shot 40.84% 70.50% 59.86% 26.00% 49.30%
Siamese[12] one-shot 70.40% 82.87% 60.62% 38.95% 63.21%
AutoFi (MLP-based)[44] one-shot 24.62% 24.71% 50.88% 23.59% 30.95%
AutoFi (CNN-based)[44] one-shot 27.05% 36.14% 48.05% 26.95% 34.55%
Yang et al.[24] one-shot 67.21% 74.22% 59.75% 48.52% 62.43%
Ding et al.[45] one-shot 39.14% 70.94% 61.56% 30.37% 50.50%
CrossFi[16] one-shot 91.72% 93.01% 80.93% 49.62% 78.82%
KNN [30] one-shot 83.02% 82.67% 49.63% 46.87% 65.55%
KNN-MMD (Ours) one-shot 93.26% 81.84% 77.62% 75.30% 82.01%
Ablation Study one-shot 69.87% 73.78% 84.03% 74.06% 75.44%
MMD[17] zero-shot 47.92% 67.25% 74.32% 45.61% 58.75%
MK-MMD[18] zero-shot 40.36% 66.47% 72.26% 43.72% 55.70%
DANN[35] zero-shot 41.41% 67.18% 74.06% 35.99% 54.66%
ADDA[46] zero-shot 42.71% 65.43% 62.81% 36.08% 51.76%
GFK+KNN[33] zero-shot 30.79% 51.50% 53.72% 34.17% 42.55%
CrossFi[16] zero-shot 64.81% 72.79% 74.38% 40.46% 63.11%
Tian et al.[41] zero-shot 68.13% 55.86% 61.72% 42.10% 56.95%
EEG[42] zero-shot 59.75% 64.63% 69.53% 42.15% 59.02%

### v2

# KNN-MMD: Cross Domain Wireless Sensing via Local Distribution Alignment

@@ Abstract
Wireless sensing has recently found widespread applications in diverse environments, including homes, offices, and public spaces. By analyzing patterns in channel state information (CSI), it is possible to infer human actions for tasks such as person identification, gesture recognition, and fall detection. However, CSI is highly sensitive to environmental changes, where even minor alterations can significantly distort the CSI patterns. This sensitivity often leads to performance degradation or outright failure when applying wireless sensing models trained in one environment to another. To address this challenge, Domain Alignment (DAL) has been widely adopted for cross-domain classification tasks, as it focuses on aligning the global distributions of the source and target domains in feature space. Despite its popularity, DAL often neglects inter-category relationships, which can lead to misalignment between categories across domains, even when global alignment is achieved. To overcome these limitations, we propose K-Nearest Neighbors Maximum Mean Discrepancy (KNN-MMD), a novel few-shot method for cross-domain wireless sensing. Our approach begins by constructing a “help set” using K-Nearest Neighbors (KNN) from the target domain, enabling local alignment between the source and target domains within each category using Maximum Mean Discrepancy (MMD). Additionally, we address a key instability issue commonly observed in cross-domain methods, where model performance fluctuates sharply between epochs. Further, most existing methods struggle to determine an optimal stopping point during training due to the absence of labeled data from the target domain. Our method resolves this by excluding the support set from the target domain during training and employing it as a validation set to determine the stopping criterion. We evaluate the effectiveness of the proposed method across several cross-domain Wi-Fi sensing tasks, including gesture recognition, person identification, fall detection, and action recognition, using both a public dataset and a self-collected dataset. In a one-shot scenario, our method achieves accuracy rates of 93.26%, 81.84%, 77.62%, and 75.30% for the respective tasks. To support future research, we will release our code and dataset to the public upon publication.
@@ I Introduction
These days, the importance of wireless sensing has grown significantly, driven by advancements in technology and the increasing demand for efficient monitoring solutions. Wireless sensing has found applications in various fields such as fall detection [1], person localization [2], and people identification [3]. Among the various manners, Wi-Fi sensing stands out as a popular method with numerous advantages, as outlined in Table I. First, Wi-Fi sensing is contactless as the object to be detected does not need to carry a device like a smartphone or smartwatch. Second, Wi-Fi sensing is privacy-preserving compared with camera. Third, Wi-Fi sensing is insensitive to occlusions and illumination conditions since its electromagnetic wave operates on a frequency band that can penetrate walls. Moreover, Wi-Fi sensing is ubiquitous, with over $19.5$ billion Wi-Fi devices deployed globally [4]. With an appropriate algorithm that enables its sensing function, ubiquitous sensing function can be realized in various environments, such as homes, offices, and public spaces. Last but not least, Wi-Fi sensing has lower cost compared to dedicated radars.
@@
Given the released tool that works with some specific types of commercial Network Interface Cards (NIC) [5], the Channel State Information (CSI) can be obtained, which describes how the wireless signals fade during the propagation from the sender to the receiver. By using this CSI data collection tool, a series of research employing commercial Wi-Fi NICs for different kinds of tasks. Moreover, a dedicated IEEE standard for Wi-Fi sensing named IEEE 802.11bf is approved [6] to enable faster development and practicalization of Wi-Fi sensing technologies. However, since the Orthogonal Frequency Division Multiplexing (OFDM) waveforms employed by Wi-Fi are designed for communication purpose without sensing optimization, it is thus challenging to apply physical models to infer and sense the ambient environment from the CSI measurements.
As a result, most Wi-Fi sensing methods rely on data-driven approach via machine learning [7, 8], especially for fine-grained tasks like gesture recognition [9]. However, even minor environmental changes can drastically alter the CSI patterns and significantly degrade the performance of the trained model. This environmental variability leads to a domain shift - a discrepancy between the training data distribution and real-world deployment conditions. Given the strong dependence of the training data, such domain shifts can severely undermine the generalization capability. This scenario is known as a cross-domain problem in machine learning, where the source domain (training data) and target domain (testing data) have different underlying distributions or feature spaces.
@@
As the most common method for cross-domain tasks, Domain Adaptation (DA) has been widely utilized in cross-domain Wi-Fi sensing. DA focuses on leveraging knowledge from a source domain to enhance model performance on a related but different target domain. Among these methods, Domain Alignment (DAL), which aims to map the source and target domains to the same distribution, is one of the most popular approaches and has been extensively applied in Wi-Fi sensing [11, 12, 13, 14, 15, 16, 17]. However, as shown in Fig. 1, the global alignment of traditional DAL methods can lead to incorrect alignments within categories, which can hinder the effectiveness of the classification boundary trained in the source domain when applied to the target domain. Another widely recognized issue in DA is low stability: the model’s performance on the testing set can be unstable during training (as referenced in [18] and Section IV-D), with accuracy varying sharply between successive epochs. This instability suggests that simply stopping the model training at a fixed epoch may result in poor performance. If we wish to implement an early stopping method, we require labeled validation data from the target domain. However, the only labeled data available from the target domain is the support set, which is often a subset of the training set and cannot be used as a validation set.
To address the limitations of traditional DAL methods in cross-domain Wi-Fi sensing, we propose a novel Few-Shot Learning (FSL) framework, called K-Nearest Neighbors Maximum Mean Discrepancy (KNN-MMD). Our approach introduces a local distribution alignment strategy that aligns the source and target domains within each category, rather than globally aligning their overall data distributions. This refinement enhances the effectiveness of traditional DAL techniques by addressing inter-category misalignments. Additionally, our framework ensures that the support set is excluded from the network’s training process, enabling the implementation of an early stopping strategy based on support set performance. This ensures stable and consistent target domain performance while avoiding the challenges faced by traditional DAL methods in determining an optimal stopping point.
The key contributions of this work are summarized as follows:
1. A Few-Shot Learning Framework with Local Alignment: We present KNN-MMD, a novel FSL framework tailored for cross-domain Wi-Fi sensing. Unlike traditional DAL methods that rely on global alignment, our framework leverages a category-specific alignment strategy to improve accuracy and robustness. Specifically, we use K-Nearest Neighbors (KNN) to assign pseudo-labels to the target domain and construct a high-confidence “help set.” This help set, in combination with the source domain training data, is then used to perform local alignment during the training of the classification network.
2. An Early Stopping Strategy Using the Support Set: Our framework introduces a practical early stopping mechanism by excluding the support set from the training process while using it as a validation set for monitoring accuracy and loss. This approach ensures the training process concludes at an optimal point, maintaining strong performance in the target domain. Traditional DAL methods often lack such a mechanism, leading to unstable or suboptimal performance.
@@
3. Comprehensive Experimental Validation: We extensively evaluate our method across multiple Wi-Fi sensing tasks, including gesture recognition, person identification, fall detection, and action recognition. Experimental results demonstrate that our framework outperforms existing models in both accuracy and stability, showcasing its effectiveness in real-world cross-domain Wi-Fi sensing scenarios.
@@
II Problem Statement and Existing Methods
@@ II-A Wi-Fi Sensing Principle
II-B Domain Shift Issue in Wi-Fi Sensing
@@
To facilitate understanding, we first define the symbols and the cross-domain scenarios. In the following, the variables $x$ and $y$ refer to the input data and output label, respectively. $\theta$ represents the network parameters. $P_{s}$ and $P_{t}$ represent the classification probabilities of each category in the source domain and target domain, respectively. Furthermore, the focus of this paper is primarily on the few-shot cross-domain task. Most available data is in the source domain expect $n$ labeled samples for each category in the target domain. This scenario is often referred to as the n-shot problem. In addition, we also have some unlabeled data from the target domain. Our objective is to accurately classify these unlabeled data. For the sake of simplicity, we will refer to the labeled data from the source domain as the training set, and the labeled data from the target domain as the support set.
In cross-domain Wi-Fi sensing, the most significant challenge is the substantial variation in data distribution across different domains, which causes models trained in the source domain to often fail in the target domain. For better illustration, we use WiGesture dataset [3], which contains the CSI collected from 8 individuals performing 6 different gesture actions, to illustrate the domain shift problem in detail. For the gesture recognition or action classification task, we take different people as different domains. And for the people identification task, we take different actions as different domains. Fig. 2 depicts the distribution of mean and standard deviation of CSI amplitude across various individuals and actions. It is evident that even for the same action, individuals exhibit distinct data distributions. Likewise, Fig. 3a and 3b display the Uniform Manifold Approximation and Projection (UMAP) based data reduction [19] results (we will introduce this data dimension reduction method in Section III-D1). It is apparent that within a single individual, different actions exhibit significant dissimilarities. However, when considering different individuals, the boundaries separating different actions become less distinct. These findings imply that it is challenging to generalize the learned features from one set of individuals to others. In Fig. 3c, a Resnet18 [20] is trained by setting the people with IDs 1-7 as the source domain and people with ID 0 as the target domain. It can be observed that the trained Resnet18 can distinguish different actions in the source domain but fail to distinguish different actions in the target domain. This is mainly because even for the same action, different people have distinct behavioral patterns, which can lead to significant gaps in the CSI patterns between individuals. As a result, to make the model perform well in the target domain, it is necessary to address the domain shift (cross-domain) problem.
@@
II-C Existing Domain Adaptation Methods
@@
Metric-based Method Learning-based Method Domain Alignment Method Ours Representative Methods KNN[21], K-means[22] Siamese[23], Triplet Network[24] MMD[25], GFK[26, 27] KNN-MMD Sensitivity to Quality of Support Set High Moderate None Low Stability Low Low Low High Assumption $P_{t}(y|x)=P_{s}(y|x)$ No Some methods require it[28] Yes No
As the most prevalent method in cross-domain tasks, DA encompasses various classification approaches. We primarily follow [29, 30], categorizing DA into three types: metric-based methods, learning-based methods, and DAL methods. In practice, a method may belong to multiple categories, however, this paper focuses on their primary classifications: metric-based methods refer mainly to traditional machine learning approaches that do not involve training, learning-based methods pertain to those associated with novel network architectures, and DAL methods primarily address the alignment of feature spaces between source and target domains.
@@
DAL methods aim to map the source domain and target domain to a common feature distribution, where the data in the two domains can have the same distribution in the new space. However, most DAL methods, such as MMD [25], Kernel Mean Matching (KMM) [34], and Subspace Geodesic Flow (SGF) [35], rely on the assumption that the conditional probabilities in the source domain and target domain are the same, i.e., $P_{t}(y|x)=P_{s}(y|x)=P(y|x)$. To estimate $P(y|x)$, the network parameters $\theta$ need to be learned to achieve $P(y|x;\theta)=P(y|x)$. However, due to $P_{t}(x)\neq P_{s}(x)$, the $P(y|x;\theta)$ learned from the source domain cannot generalize to the target domain. Consequently, these traditional methods can be seen as finding a latent space where the input data from the source domain and target domain are mapped to the same distribution, i.e., $P_{t}(x)=P_{s}(x)$. This allows the model to learn the parameter $\theta$ more effectively and accurately. However, in many practical scenarios, the assumption $P_{t}(y|x)=P_{s}(y|x)$ is not met. This has been demonstrated in our experiments (Section IV-C) as well as our previous research [18], where traditional DAL methods even failed in cross-domain Wi-Fi sensing tasks.
@@ III-C Model Structure
Our model architecture is illustrated in Fig. 6, which primarily consists of three components: a normalization layer, a feature extractor, and a final classifier. The normalization layer incorporates time-level standardization and batch normalization, helping to reduce the domain gap between different samples and facilitating the training process. Then we select the pre-trained ResNet18 [20] as the feature extractor because the Wi-Fi signal spectrum exhibits high similarity to images. Convolutional networks, which effectively extract local features, have demonstrated excellent performance in CSI classification [18, 1, 40]. Furthermore, inspired by transfer learning, the pre-trained parameters from image data provide a strong starting point for Wi-Fi sensing [41]. Finally, we employ a MLP to process the embeddings for classification. The detailed description is as follows:
@@ III-D Training Process
The training process consists of two main phases: preliminary classification and network training with local alignment. In order to align the source and target domains locally within each category, we need to know the labels of the target domain samples. However, since we have only a very limited labeled support set in the target domain, it is challenging to use these samples for alignment, as this typically requires a larger number of samples to capture statistical features effectively. To address this issue, we propose using KNN for preliminary classification of the target domain. Although KNN may not be very stable and often produces numerous incorrect classifications, we find that classification accuracy is closely related to the confidence level. By leveraging this insight, we can construct a helper set consisting of samples with high confidence levels, the majority of which are correctly classified in practice. During the network training phase, we use this helper set to achieve local alignment with the source domain training set. Finally, the trained network can be employed for final classification, demonstrating excellent performance on samples with low classification confidence from KNN. Moreover, since the support set does not participate in network training, we can utilize it as a validation set to determine when to stop model training. In the following, we provide a detailed description of the two phases:
@@ IV Experiment
Figure 6: Hardware: Our system utilizes a commercial Wi-Fi router as the transmitter (TX) and an ESP32-S3 as the receiver (RX). The ESP32-S3 decodes the received Channel State Information (CSI) and transmits it to a PC for inference.
@@ IV-A Dataset Description
After training, the model can be integrated into a real-time system, as shown in Fig. 6. In this system, a commercial Wi-Fi router continuously sends ping reply packets to the ESP32-S3. The ESP32-S3 extracts the CSI and forwards it to a PC equipped with an NVIDIA RTX 3090 GPU. The PC can then perform real-time inference using the trained neural network.
@@ IV-C Experiment Result
Method Scenario Gesture Recognition People Identification Fall Detection Action Recognition Average Accuracy Resnet18[20] in-domain 80.75% 86.75% 91.88% 70.50% 82.47% zero-shot 40.84% 70.50% 59.86% 26.00% 49.30% Siamese[23] one-shot 70.40% 82.87% 60.62% 38.95% 63.21% AutoFi (MLP-based)[43] one-shot 24.62% 24.71% 50.88% 23.59% 30.95% AutoFi (CNN-based)[43] one-shot 27.05% 36.14% 48.05% 26.95% 34.55% Yang et al.[16] one-shot 67.21% 74.22% 59.75% 48.52% 62.43% Ding et al.[44] one-shot 39.14% 70.94% 61.56% 30.37% 50.50% CrossFi[18] one-shot 91.72% 93.01% 80.93% 49.62% 78.82% KNN [21] one-shot 83.02% 82.67% 49.63% 46.87% 65.55% KNN-MMD (Ours) one-shot 93.26% 81.84% 77.62% 75.30% 82.01% Ablation Study one-shot 69.87% 73.78% 84.03% 74.06% 75.44% MMD[25] zero-shot 47.92% 67.25% 74.32% 45.61% 58.75% MK-MMD[39] zero-shot 40.36% 66.47% 72.26% 43.72% 55.70% DANN[28] zero-shot 41.41% 67.18% 74.06% 35.99% 54.66% ADDA[45] zero-shot 42.71% 65.43% 62.81% 36.08% 51.76% GFK+KNN[26] zero-shot 30.79% 51.50% 53.72% 34.17% 42.55% CrossFi[18] zero-shot 64.81% 72.79% 74.38% 40.46% 63.11% Tian et al.[36] zero-shot 68.13% 55.86% 61.72% 42.10% 56.95% EEG[37] zero-shot 59.75% 64.63% 69.53% 42.15% 59.02%